a block of 18kg is pushed along a horizontal frictionless surface by a horizontal force of 80N. the block start from rest. compute the kinetic energy after 6 seconds


1
Expert's answer
2022-03-27T10:56:04-0400

"a=\frac Fm,"

"v=at=\frac{Ft}m,"

"E=\frac{mv^2}2=\frac{F^2t^2}{2m}=6.4~kJ."